Problem 9776: VELO

System
VELO
Severity
Report
Start date
2017-05-28 11:09
End date
2017-06-20 17:04
Runs affected
Reported by
Wenhua Hu on 2017-05-28 11:09

Comments

#1 by Wenhua Hu on 2017-05-28 11:09

At least one VELO sensor has 0% occupancy.
Shift leader had tried fast run chan1ge, the colour of plot truns to be white.
I'm not sure if it is a problem.

Logbook link

#2 by Kurt Rinnert on 2017-06-20 17:04

This was due to a wrong configuration of an updated version of the monitoring algorithm. A fall back to the previous version fixed the problem.
Closed the problem with end date: 2017-06-20 17:04:00

You can sign in and add comments to this problem.